apmar:

Using the following equation, find the center in radius of the circle. You must show all work and calculations to receive credit. X^2+2x+y^2+4y=20

L1ly:

x^2+2x+y^2+4y=20 (x^2 + 2x) + (y^2 + 4y) = 20 Using the completing the square method: (x^2 + 2x + 1) + (y^2 + 4y + 4) = 20 + 1 + 4 (x + 1)^2 + (y + 2)^2 = 25 (x + 1)^2 + (y + 2)^2 = 5^2 This is in the format (x-h)^2 + (y-k)^2 = r^2, where (h,k) are the center of the circle and r is the radius. Can you figure out what the center and radius are now?